# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 2.6. Trivial algebras
Since ${\mathcal{D}}$ is bitensored over monoidal $\infty$-categories, by [19, Lemma 3.103.] the left adjoint ${\mathrm{L}}$ of the embedding ${\mathcal{D}}\subset{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$ preserves the biactions and so induces a localization ${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{L}}):{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\rightleftarrows{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})$. Thus ${\mathrm{Bar}}^{{\mathcal{D}}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})$ factors as
${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})\subset{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\xrightarrow{{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{L}})}{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})\xrightarrow{{\mathrm{Bar}}^{{\mathcal{D}}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathcal{D}})$
and ${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\xrightarrow{{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{L}})}{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})\xrightarrow{{\mathrm{Bar}}^{{\mathcal{D}}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathcal{D}})$ identifies with
${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\xrightarrow{{\mathrm{Bar}}^{{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\xrightarrow{\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathrm{L}})}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathcal{D}}).$
@@
Let ${\mathcal{D}}$ be an $\infty$-category lax bitensored over non-symmetric $\infty$-operads ${\mathcal{C}},{\mathcal{E}}$ and ${\mathrm{A}}$ an algebra in ${\mathcal{C}}$, ${\mathrm{B}}$ an algebra in ${\mathcal{E}}$. For any right ${\mathrm{A}}$-module ${\mathrm{X}}$ and left ${\mathrm{B}}$-module ${\mathrm{Y}}$ let ${\mathrm{Bar}}^{\mathcal{D}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})$ be the composition:
${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})\subset{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\xrightarrow{{\mathrm{Bar}}^{{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))$
@@
For any ${\mathrm{M}}\in{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})$ and $[{\mathrm{n}}]\in\Delta$ we have
${\mathrm{Bar}}^{\mathcal{D}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},{\mathrm{M}},{\mathrm{Y}})_{\mathrm{n}}\simeq{\mathrm{X}}\otimes\underbrace{{\mathrm{A}}\otimes...\otimes{\mathrm{A}}}_{{\mathrm{n}}-\text{times}}\otimes{\mathrm{M}}\otimes\underbrace{{\mathrm{B}}\otimes...\otimes{\mathrm{B}}}_{{\mathrm{n}}-\text{times}}\otimes{\mathrm{Y}}.$
@@
For every ${\mathrm{A}},{\mathrm{B}}$-bimodule ${\mathrm{M}}$ let ${\mathrm{X}}\otimes_{\mathrm{A}}{\mathrm{M}}\otimes_{\mathrm{B}}{\mathrm{Y}}$ be the geometric realization of ${\mathrm{Bar}}^{\mathcal{D}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},{\mathrm{M}},{\mathrm{Y}})$ if the latter exists.
@@
Let ${\mathcal{D}}$ be an $\infty$-category lax bitensored over non-symmetric $\infty$-operads ${\mathcal{C}},{\mathcal{E}}$, ${\mathrm{A}}$ an augmented algebra in ${\mathcal{C}}$ and ${\mathrm{B}}$ an augmented algebra in ${\mathcal{E}}$. If ${\mathcal{D}}$ has geometric realizations, the functor
@@
By [23, Remark 4.5.3.3.] there is an adjunction
@@
Since the biaction of ${\mathcal{C}},{\mathcal{E}}$ on ${\mathcal{D}}$ is lax, the forgetful functor ${{}_{\mathbb{1}}}{\mathrm{BMod}}_{\mathbb{1}}({\mathcal{D}})\to{\mathcal{D}}$ is an equivalence. Let ${\mathrm{L}}$ be the left adjoint of the embedding ${\mathcal{D}}\subset{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$. When viewing ${\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$ as bitensored over the full monoidal subcategories of ${\mathrm{Env}}({\mathcal{C}}),{\mathrm{Env}}({\mathcal{E}})$ spanned by the tensor unit of ${\mathcal{C}},{\mathcal{E}}$, respectively, and ${\mathcal{D}}$ as bitensored over the full subcategories of ${\mathcal{C}},{\mathcal{E}}$ spanned by the tensor unit of ${\mathcal{C}},{\mathcal{E}}$, respectively, the functor ${\mathrm{L}}:{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})\to{\mathcal{D}}$ preserves the biaction and so gives rise to a localization
@@
${{}_{\mathbb{1}}}{\mathrm{BMod}}_{\mathbb{1}}({\mathcal{P}}{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))(\mathrm{colim}({\mathrm{Bar}}_{{\mathrm{A}},{\mathrm{B}}}^{{\mathcal{P}}{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})}({\mathbb{1}},{\mathrm{M}},{\mathbb{1}})),{\mathrm{Y}})\simeq$
$\lim{{}_{\mathbb{1}}}{\mathrm{BMod}}_{\mathbb{1}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))({\mathrm{Bar}}^{{\mathcal{B}}{\mathrm{Env}}({\mathcal{D}})}_{{\mathrm{A}},{\mathrm{B}}}({\mathbb{1}},{\mathrm{M}},{\mathbb{1}}),{\mathrm{Y}})\simeq$
$\lim{\mathcal{D}}({\mathrm{L}}\circ{\mathrm{Bar}}^{{\mathcal{B}}{\mathrm{Env}}({\mathcal{D}})}_{{\mathrm{A}},{\mathrm{B}}}({\mathbb{1}},{\mathrm{M}},{\mathbb{1}}),{\mathrm{Y}})\simeq{\mathcal{D}}(\mathrm{colim}({\mathrm{L}}\circ{\mathrm{Bar}}^{{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})}_{{\mathrm{A}},{\mathrm{B}}}({\mathbb{1}},{\mathrm{M}},{\mathbb{1}})),{\mathrm{Y}}).$
3. Restricted $L_{\infty}$-algebras
In this section we define the monad of tensor algebra / primitive elements whose algebras we call restricted $L_{\infty}$-algebras.
Construction 3.1.
Let ${\mathcal{C}}$ be an additive presentably symmetric monoidal $\infty$-category.
The symmetric monoidal $\infty$-category ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ is compatible with small colimits and the forgetful functor ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ preserves small colimits. This implies that the forgetful functor
$\mathrm{Bialg}({\mathcal{C}})={\mathrm{Grp}}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))=\mathrm{Mon}({\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))\to{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
admits a left adjoint $\widetilde{\mathrm{Free}}$ that covers the left adjoint of the forgetful functor $\mathrm{Alg}({\mathcal{C}})\to{\mathcal{C}}_{{\mathbb{1}}/},$ the free associative algebra functor.
By Corollaries 2.54 and 2.57 there is an equivalence
$\mathrm{Coalg}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\simeq{\mathrm{coAlg}}_{{\mathrm{coComm}}}({\mathcal{C}})_{{\mathbb{1}}/}\simeq{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
@@ 3. Restricted $L_{\infty}$-algebras
${\mathcal{C}}\xrightarrow{{\mathrm{triv}}_{{\mathrm{coComm}}^{\mathrm{nu}}}}{\mathrm{coAlg}}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\simeq{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}\xrightarrow{\widetilde{\mathrm{Free}}}\mathrm{Bialg}({\mathcal{C}}).$
@@
$\mathrm{Bialg}({\mathcal{C}})\xrightarrow{\mathrm{forget}}{\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}\simeq{\mathrm{coAlg}}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\xrightarrow{{\mathrm{triv}}*^{{\mathrm{coComm}}^{\mathrm{nu}}}(-)}{\mathcal{C}}.$
Then there is an adjunction
${\mathrm{T}}:{\mathcal{C}}\rightleftarrows\mathrm{Bialg}({\mathcal{C}}):{\mathrm{Prim}}.$
Lemma 3.2.
Let ${\mathcal{C}}$ be an additive presentably symmetric monoidal $\infty$-category. The tensor algebra functor
${\mathrm{T}}:{\mathcal{C}}\to\mathrm{Bialg}({\mathcal{C}})\simeq\mathrm{Mon}({\mathrm{Cocoalg}}({\mathcal{C}}))$
induces a functor ${\mathcal{C}}\to{\mathrm{Hopf}}({\mathcal{C}})={\mathrm{Grp}}({\mathrm{Cocoalg}}({\mathcal{C}}))\subset\mathrm{Mon}({\mathrm{Cocoalg}}({\mathcal{C}}))$.
Proof.
Let ${\mathrm{X}}\in{\mathcal{C}}$ and $\mathrm{pr}_{1},\mathrm{pr}_{2}:{\mathrm{X}}\times{\mathrm{X}}\to{\mathrm{X}}$ the projections and $\mu:{\mathrm{X}}\times{\mathrm{X}}\simeq{\mathrm{X}}\coprod{\mathrm{X}}\to{\mathrm{X}}$ the codiagonal. Since ${\mathrm{X}}\in{\mathcal{C}}\simeq{\mathrm{Grp}}({\mathcal{C}})$ is a group object, the canonical morphism $\alpha:{\mathrm{X}}\times{\mathrm{X}}\xrightarrow{(\mathrm{pr}_{1},\mu)}{\mathrm{X}}\times{\mathrm{X}}$ is an equivalence. We like to see that ${\mathrm{T}}({\mathrm{X}})\in\mathrm{Bialg}({\mathcal{C}})$ is a Hopf algebra. Since the canonical functor $\mathrm{Bialg}({\mathcal{C}})\to\mathrm{Bialg}(\mathsf{Ho}({\mathcal{C}}))$ is conservative and preserves finite products, i.e. the tensor product, it detects Hopf algebras. So it is enough to see that the image of ${\mathrm{T}}({\mathrm{X}})$ in the 1-category $\mathrm{Bialg}(\mathsf{Ho}({\mathcal{C}}))$ is a Hopf algebra. For that by classical theory it is enough to check that ${\mathrm{T}}({\mathrm{X}})$ admits an antipode. Consequently, we can assume that ${\mathcal{C}}$ is a 1-category. Let ${\mathrm{i}}:{\mathrm{X}}\simeq{\mathrm{X}}\times 0\to{\mathrm{X}}\times{\mathrm{X}}\xrightarrow{\alpha^{-1}}{\mathrm{X}}\times{\mathrm{X}}\xrightarrow{\mathrm{pr}_{2}}{\mathrm{X}}$ be the inverse of ${\mathrm{X}}.$ The morphism ${\mathrm{j}}:={\mathrm{T}}({\mathrm{i}}):{\mathrm{T}}({\mathrm{X}})\to{\mathrm{T}}({\mathrm{X}})$ is an antipode for ${\mathrm{A}}:={\mathrm{T}}({\mathrm{X}})$ since the commutative squares
in ${\mathcal{C}}$ give rise to commutative diagrams
in ${\mathcal{C}}$, where $\ast$ is the coproduct in $\mathrm{Alg}({\mathcal{C}}).$
Corollary 3.3.
Let ${\mathcal{C}}$ be an additive presentably symmetric monoidal $\infty$-category. The adjunction
restricts to an adjunction
${\mathrm{T}}:{\mathcal{C}}\rightleftarrows{\mathrm{Hopf}}({\mathcal{C}}):{\mathrm{Prim}}$
denoted by the same name.
Definition 3.4.
Let ${\mathcal{C}}$ be an additive presentably symmetric monoidal $\infty$-category.
A restricted $L_{\infty}$-algebra is an algebra for the monad associated to the adjunction
${\mathrm{T}}:{\mathcal{C}}\rightleftarrows{\mathrm{Hopf}}({\mathcal{C}}):{\mathrm{Prim}}.$
Example 3.5.
Restricted $L_{\infty}$-algebras in the category of (differentially) (graded) ${\mathrm{K}}$-vector spaces agree with restricted (differentially) (graded) Lie algebra over ${\mathrm{K}}$. This holds by a result of Fresse [10, Theorem 1.2.5.] or likewise the version of the Milnor-Moore theorem in positive characteristic [24, Theorem 5.18.].

### v2

@@ 2.6. Trivial algebras
Since $\mathcal{D}$} is bitensored over monoidal $\infty$-categories, by [25, Lemma 3.103.] the left adjoint $\mathrm{L}$} of the embedding ${\mathcal{D}}\subset{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$ preserves the biactions and so induces a localization ${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{L}}):{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\rightleftarrows{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})$. Thus ${\mathrm{Barc}}^{{\mathcal{D}}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})$ factors as
${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})\subset{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\xrightarrow{{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{L}})}{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})\xrightarrow{{\mathrm{Barc}}^{{\mathcal{D}}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathcal{D}})$
and ${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\xrightarrow{{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{L}})}{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})\xrightarrow{{\mathrm{Barc}}^{{\mathcal{D}}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathcal{D}})$ identifies with
${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\xrightarrow{{\mathrm{Barc}}^{{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\xrightarrow{\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathrm{L}})}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathcal{D}}).$
@@
Let $\mathcal{D}$} be an $\infty$-category lax bitensored over non-symmetric $\infty$-operads $\mathcal{C}$$\mathcal{E}$} and $\mathrm{A}$} an algebra in $\mathcal{C}$}, $\mathrm{B}$} an algebra in $\mathcal{E}$}. For any right $\mathrm{A}$}-module $\mathrm{X}$} and left $\mathrm{B}$}-module $\mathrm{Y}$} let ${\mathrm{Barc}}^{\mathcal{D}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})$ be the composition:
${{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})\subset{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))\xrightarrow{{\mathrm{Barc}}^{{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},-,{\mathrm{Y}})}\mathrm{Fun}(\Delta^{\mathrm{op}},{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))$
@@
For any ${\mathrm{M}}\in{{}_{\mathrm{A}}}{\mathrm{BMod}}_{\mathrm{B}}({\mathcal{D}})$ and ${n}]\in\Delta$ we have
${\mathrm{Barc}}^{\mathcal{D}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},{\mathrm{M}},{\mathrm{Y}})_{n}\simeq{\mathrm{X}}\otimes\underbrace{{\mathrm{A}}\otimes...\otimes{\mathrm{A}}}_{{n}-\text{times}}\otimes{\mathrm{M}}\otimes\underbrace{{\mathrm{B}}\otimes...\otimes{\mathrm{B}}}_{{n}-\text{times}}\otimes{\mathrm{Y}}.$
@@
For every ${\mathrm{A}},{\mathrm{B}}$-bimodule $\mathrm{M}$} let ${\mathrm{X}}\otimes_{\mathrm{A}}{\mathrm{M}}\otimes_{\mathrm{B}}{\mathrm{Y}}$ be the geometric realization of ${\mathrm{Barc}}^{\mathcal{D}}_{{\mathrm{A}},{\mathrm{B}}}({\mathrm{X}},{\mathrm{M}},{\mathrm{Y}})$ if the latter exists.
@@
Let $\mathcal{D}$} be an $\infty$-category lax bitensored over non-symmetric $\infty$-operads $\mathcal{C}$$\mathcal{E}$}, $\mathrm{A}$} an augmented algebra in $\mathcal{C}$} and $\mathrm{B}$} an augmented algebra in $\mathcal{E}$}. If $\mathcal{D}$} has geometric realizations, the functor
@@
By [39, Remark 4.5.3.3.] there is an adjunction
@@
Since the biaction of $\mathcal{C}$$\mathcal{E}$} on $\mathcal{D}$} is lax, the forgetful functor ${{}_{\mathbb{1}}}{\mathrm{BMod}}_{\mathbb{1}}({\mathcal{D}})\to{\mathcal{D}}$ is an equivalence. Let $\mathrm{L}$} be the left adjoint of the embedding ${\mathcal{D}}\subset{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$. When viewing ${\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})$ as bitensored over the full monoidal subcategories of ${\mathrm{Env}}({\mathcal{C}}),{\mathrm{Env}}({\mathcal{E}})$ spanned by the tensor unit of $\mathcal{C}$$\mathcal{E}$}, respectively, and $\mathcal{D}$} as bitensored over the full subcategories of $\mathcal{C}$$\mathcal{E}$} spanned by the tensor unit of $\mathcal{C}$$\mathcal{E}$}, respectively, the functor ${\mathrm{L}}:{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})\to{\mathcal{D}}$ preserves the biaction and so gives rise to a localization
@@
${{}_{\mathbb{1}}}{\mathrm{BMod}}_{\mathbb{1}}({\mathcal{P}}{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))(\mathrm{colim}({\mathrm{Barc}}_{{\mathrm{A}},{\mathrm{B}}}^{{\mathcal{P}}{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})}({\mathbb{1}},{\mathrm{M}},{\mathbb{1}})),{\mathrm{Y}})\simeq$
$\lim{{}_{\mathbb{1}}}{\mathrm{BMod}}_{\mathbb{1}}({\mathrm{B}}{\mathrm{Env}}({\mathcal{D}}))({\mathrm{Barc}}^{{\mathcal{B}}{\mathrm{Env}}({\mathcal{D}})}_{{\mathrm{A}},{\mathrm{B}}}({\mathbb{1}},{\mathrm{M}},{\mathbb{1}}),{\mathrm{Y}})\simeq$
$\lim{\mathcal{D}}({\mathrm{L}}\circ{\mathrm{Barc}}^{{\mathcal{B}}{\mathrm{Env}}({\mathcal{D}})}_{{\mathrm{A}},{\mathrm{B}}}({\mathbb{1}},{\mathrm{M}},{\mathbb{1}}),{\mathrm{Y}})\simeq{\mathcal{D}}(\mathrm{colim}({\mathrm{L}}\circ{\mathrm{Barc}}^{{\mathrm{B}}{\mathrm{Env}}({\mathcal{D}})}_{{\mathrm{A}},{\mathrm{B}}}({\mathbb{1}},{\mathrm{M}},{\mathbb{1}})),{\mathrm{Y}}).$
Next we investigate the relationship between the free associative algebra and the primitives:
Lemma 2.70.
Let $\mathcal{C}$} be an additive presentably symmetric monoidal $\infty$-category. There is an adjunction
whose left adjoint lifts the free associative algebra.
Proof.
The symmetric monoidal $\infty$-category $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})$ is compatible with small colimits and the forgetful functor $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})\to{\mathcal{C}}$ preserves small colimits. This implies that the forgetful functor
${\mathrm{Bialg}}({\mathcal{C}})={\mathrm{Grp}}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))=\mathrm{Mon}(\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}}))\to\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
admits a left adjoint $\widetilde{\mathrm{Free}}$ that covers the left adjoint of the forgetful functor $\mathrm{Alg}({\mathcal{C}})\to{\mathcal{C}}_{{\mathbb{1}}/},$ the free associative algebra functor. By Corollaries 2.54 and 2.57 there is an equivalence
$\mathrm{Coalg}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\simeq\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}}({\mathcal{C}})_{{\mathbb{1}}/}\simeq\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}$
@@
${\mathcal{C}}\xrightarrow{{\mathrm{triv}}_{{\mathrm{coComm}}^{\mathrm{nu}}}}\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\simeq\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}\xrightarrow{\widetilde{\mathrm{Free}}}{\mathrm{Bialg}}({\mathcal{C}}).$
@@
${\mathrm{Bialg}}({\mathcal{C}})\xrightarrow{\mathrm{forget}}\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{C}})_{{\mathbb{1}}/}\simeq\mathrm{co}\mathrm{Alg}_{{\mathrm{coComm}}^{\mathrm{nu}}}({\mathcal{C}})\xrightarrow{{\mathrm{triv}}*^{{\mathrm{coComm}}^{\mathrm{nu}}}(-)}{\mathcal{C}}.$
Hence there is an adjunction ${\mathrm{T}}:{\mathcal{C}}\rightleftarrows{\mathrm{Bialg}}({\mathcal{C}}):{\mathrm{Prim}}.$